28x-2x(10x-0.365)=7x+0.035
Please, please explain!

Answer by stanbon(75887)

28x-2x(10x-0.365)=7x+0.035
-----------
28x-20x^2+0.730x = 7x + 0.035
28.73x - 20x^2 = 7x + 0.035
20x^2 - 21.73x + 0.035 = 0
-------
x = [21.73 +- sqrt(21.73^2 -4*20*0.035)]/40
x = [21.73 +- sqrt(469.3920)]/40
x = [21.73 +- 21.664377...]/40
x = 1.0848869... or x = 0.0016131...
=========================
Cheers,
Stan H.